# Random Forest for Dynamic Risk Prediction of Recurrent Events: A Pseudo-Observation Approach

### v2

@@ Abstract
Recurrent events are common in clinical, healthcare, social and behavioral studies. To overcome some long-standing challenges in analyzing censored recurrent event data, a recent regression analysis framework constructs a censored longitudinal data set consisting of times to the first recurrent event in multiple pre-specified follow-up windows of length $\tau$ (Xia, Murray and Tayob, 2020, XMT). Traditional regression models struggle with nonlinear and multi-way interactions, with success depending on the skill of the statistical programmer. With a staggering number of potential predictors being generated from genetic, -omic, and electronic health records sources, machine learning approaches such as the random forest regression are growing in popularity, as they can nonparametrically incorporate information from many predictors with nonlinear and multi-way interactions involved in prediction. In this paper, we (1) develop a random forest approach for dynamically predicting probabilities of remaining event-free during a subsequent $\tau$-duration follow-up period from a reconstructed censored longitudinal data set, (2) modify the XMT regression approach to predict these same probabilities, subject to the limitations that traditional regression models typically have, and (3) demonstrate how to incorporate patient-specific history of recurrent events for prediction in settings where this information may be partially missing. We show the increased ability of our random forest algorithm for predicting the probability of remaining event-free over a $\tau$-duration follow-up period when compared to our modified XMT method for prediction in settings where association between predictors and recurrent event outcomes is complex in nature. We also show the importance of incorporating past recurrent event history in prediction algorithms when event times are correlated within a subject. The proposed random forest algorithm is demonstrated using recurrent exacerbation data from the Azithromycin for the Prevention of Exacerbations of Chronic Obstructive Pulmonary Disease (Albert and others, 2011). Censored data; Longitudinal data; Pseudo-observations; Random forest; Recurrent events
@@ 1 Introduction
Recurrent events are common in healthcare settings, such as exacerbations of pulmonary diseases, or bleeding events for chemotherapy patients. Clinicians are often interested in predicting the probability of remaining event-free until a patient’s next check-up. Information from events over time helps improve precision of predictions relative to analyses of a single time-to-event. Prediction is also greatly enhanced by the large number of patient-specific predictors that are collected now, for instance, electronic health records data, gene expression levels, metabolomic data, etc. The very large dimensions of the predictors and the multivariate nature of recurrent event outcomes requires careful thought for analysis, particularly when censoring of recurrent events is a feature of the data. A method that can handle both highly correlated data and recurrent events would be of interest to modelers and practitioners.
@@
The majority of literature in recurrent event analysis has been parametric or semi-parametric in nature, with specific modeled relationships between predictors and recurrent event outcomes. Models have been designed for the analysis of recurrent events, gap times, and times-to-first-event across different follow-up windows, each with different assumptions imposed by the modeling paradigms (e.g., Andersen and Gill, 1982; Miloslavsky and others, 2004; Ang and others, 2009; Mao and Lin, 2016; Hengelbrock and others, 2016; Kong and others, 2016; Chan and Wang, 2017a; Liang and others, 2017; Xu and others, 2017; Choi and others, 2017; Chan and Wang, 2017b; Xia, Murray and Tayob, 2020; Safari and others, 2023). Such models have interpretable parameter estimates that can inform medical literature, generate hypotheses, and to some degree, predict outcomes. Disadvantages include limits to estimating these same parameters when there are many predictors or co-linearity between important predictors. In addition, methods that incorrectly assume independence of gap times between recurrent events measured on an individual give biased results due to induced dependent censoring. A few of these papers use pseudo-observation approaches to model recurrent event data (Yokota and Matsuyama, 2019; Xia, Murray and Tayob, 2020), an attractive approach that we also consider.
@@
As to the goal of modeling many potentially collinear predictors, perhaps with non-linear associations with recurrent event outcomes, we are drawn to the nonparametric nature of random forest algorithms. Machine learning methods, and in particular random forest algorithms, have gained widespread acceptance as a method of building predictions based on complex relationships of biomarkers (first proposed by Breiman and others (1984), applied more recently in statistical literature such as Ishwaran and others (2004), Wager and Athey (2018), Zhao and others (2020), Gerber and others (2021a), Gerber and others (2021b), and reviewed in Hu and Szymczak (2023)). These algorithms naturally accommodate non-canonical relationships between covariates and outcomes, an attractive feature in an age of increasingly high-dimensional data where a traditional modeler cannot realistically find all the interactions, higher-order covariates and non-linear patterns that may be predictive of an outcome. By using information across different sets of covariates within each of the generated trees of a random forest, a modeler can side-step several traditional pitfalls of parametric and semi-parametric models: (1) multicolinearity of predictors and (2) limited ability to include a large number of covariates. In particular, Zhao and others (2020) showed that for a single event time, imposing a censored longitudinal data structure introduced by Tayob and Murray (2015) and incorporating ideas from pseudo-observation literature, random forest regression can address nonlinear relationships and interactions between covariates and outcomes for dynamic prediction of event-free periods.
In our manuscript we incorporate ideas from Xia, Murray and Tayob (2020), who propose a censored longitudinal data structure for the analysis of recurrent event data. The Xia, Murray and Tayob (XMT) model estimates restricted mean time-to-next-recurrent-event based on clinical characteristics. As part of this manuscript, we will develop a modified version of the XMT model that instead estimates event-free probabilities over $\tau$-duration follow-up windows. This more standard regression approach to prediction will be compared to a novel random forest dynamic prediction model for potentially censored recurrent event data that we introduce in this manuscript and recommend when prediction is the primary goal of analysis and when many covariates are available. For both of these modeling approaches, we explore the importance of patient-specific historical event-rate as a predictor. To date, we are not aware of any methods for dynamic prediction applicable to censored recurrent event data.
@@
The rest of this paper is organized as follows. In Section 2, we describe how to transform recurrent event data to a censored longitudinal data framework for analysis using our proposed methods. In Section 3, we review a random forest algorithm applicable to uncensored longitudinal data. In Section 4, we describe how to construct pseudo-observations based on the censored longitudinal data structure, described in Section 2, that can be used with the random forest algorithm developed in Section 3. In Section 5, we compare via simulation our modified XMT regression approach versus our proposed random forest approach for the analysis of censored recurrent event data, and evaluate different methods for including historical event rate information in these analyses. We apply both the modified XMT and our recurrent event random forest algorithm to data from Azithromycin for the Prevention of Exacerbation of COPD (Albert and others, 2011) in Section 6. We conclude with a discussion in Section 7.
@@ 2 Notation and Censored Longitudinal Data Structure
For individual $i$, denote recurrent event times $T_{i,1}<T_{i,2}<\ldots$, that are potentially censored by the random variable $C_{i}$, where $C_{i}$ is independent of $T_{i,j}$, $i=1,...,n$, and $j\geq 1$. Observed data for each individual is $X^{*}_{i,j}=\min(T_{i,j},C_{i})$, $i=1,\ldots,n$, and $j\geq 1$. These event times are sometimes converted to gap times between recurrent events $G_{i,j}=T_{i,j+1}-T_{i,j}$, $i=1,2,..n$, $j\geq 1$. A well-known challenge in the analysis of correlated censored gap times is the dependence between gap time random variables, $G_{i,j}$, and corresponding censoring times, $C_{i}-G_{i,j-1}$ (Lin and others, 2000).
One approach to avoiding dependent censoring in analyses of recurrent events was proposed by Tayob and Murray (2015), who transformed correlated recurrent event times into a censored longitudinal data structure. First a series of regularly spaced follow-up window start times are spaced every $a$ units apart, so that follow-up window start times are $\mathcal{T}=\{t_{0},t_{1},t_{2},\ldots\}$, with $t_{j}-t_{j-1}=a$, $j=1,2,\ldots$. Briefly, $\mathcal{T}$ can include window start times that are as close as $a=1$ day apart, if desired. Computational speed of the analysis is usually the limiting factor in choosing $a$. Xia and Murray (2019) found that $a$ equal to one-third of the mean gap time in the control group resulted in capturing 90% of the recurrent events from the original data set in at least one follow-up window. In practice, more frequent follow-up window start times have the potential for more precise estimation, but with diminishing returns once all recurrent events have been captured in at least one follow up window.
In this paper, at each of the pre-specified “check-in time points” in $\mathcal{T}$, we will focus on estimating the probability of remaining recurrent-event-free over a subsequent follow-up window of length $\tau$, where $\tau$ is user-specified and ideally informed by clinical expertise. Although the check-in time points are shared across subjects by design, potential censoring may result in different numbers of effective check-ins; we therefore denote the subject-specific sets of check-in time points by $\{\mathcal{T}_{i}\}_{i=1}^{n}$, where $\mathcal{T}_{i}=\{t_{0},t_{1},...t_{m_{i}}\}$, $t_{m_{i}}<\min\{C_{i},D-\tau\}$, and $D$ is the time of administrative censoring, e.g., indicating the end of study.
We follow Tayob and Murray (2015) and construct the censored longitudinal data that is essential for creating pseudo-observations in Section 4. We repeat the following two steps at all the check-in time points $t\in\mathcal{T}_{i}$, $i=1,\ldots,n$:
@@
    Identify the first recurrent event occurred after $t$ and let $\eta_{i}(t)=\min\{j:X^{*}_{i,j}\geq t,j=1,\ldots,n_{i}\}$ be its index among all the observed recurrent events for subject $i$; also define the residual recurrent-event-free time by $T_{i}(t)=T_{i,\eta_{i}(t)}-t$ with the corresponding censoring random variable $C_{i}(t)=C_{i}-t$, which remains independent of $T_{i}(t)$;
@@
    Construct the observed censored longitudinal data set at time $t$ by $\{X_{i}(t)=X^{*}_{i,\eta_{i}(t)}-t,\delta_{i}(t)=I(X_{i}(t)\leq C_{i}(t)):i=1,2,...n\}$.

### v3

@@ Abstract
Recurrent events are common in clinical, healthcare, social and behavioral studies, yet methods for dynamic risk prediction of these events are limited. To overcome some long-standing challenges in analyzing censored recurrent event data, a recent regression analysis framework constructs a censored longitudinal data set consisting of times to the first recurrent event in multiple pre-specified follow-up windows of length $\tau$ (Xia, Murray and Tayob, 2020, XMT). Traditional regression models struggle with nonlinear and multi-way interactions, with success depending on the skill of the statistical programmer. With a staggering number of potential predictors being generated from genetic, -omic, and electronic health records sources, machine learning approaches such as the random forest regression are growing in popularity, as they can nonparametrically incorporate information from many predictors with nonlinear and multi-way interactions involved in prediction. In this paper, we (1) develop a random forest approach for dynamically predicting probabilities of remaining event-free during a subsequent $\tau$-duration follow-up period from a reconstructed censored longitudinal data set, (2) modify the XMT regression approach to predict these same probabilities, subject to the limitations that traditional regression models typically have, and (3) demonstrate how to incorporate patient-specific history of recurrent events for prediction in settings where this information may be partially missing. We show the increased ability of our random forest algorithm for predicting the probability of remaining event-free over a $\tau$-duration follow-up window when compared to our modified XMT method for prediction in settings where association between predictors and recurrent event outcomes is complex in nature. We also show the importance of incorporating past recurrent event history in prediction algorithms when event times are correlated within a subject. The proposed random forest algorithm is demonstrated using recurrent exacerbation data from the trial of Azithromycin for the Prevention of Exacerbations of Chronic Obstructive Pulmonary Disease (Albert and others, 2011). Censored data; Longitudinal data; Pseudo-observations; Random forest; Recurrent events
@@ 1 Introduction
Recurrent events are common in healthcare settings, such as exacerbations of pulmonary diseases, or bleeding events for chemotherapy patients. Clinicians are often interested in predicting the probability of remaining event-free until a patient’s next check-up. Information from events over time helps improve precision of predictions relative to analyses of a single time-to-event. Prediction is also greatly enhanced by the large number of patient-specific predictors that are collected now, for instance, in the form of electronic health records data, gene expression levels, metabolomic data, etc. The very large dimensions of the predictors and the multivariate nature of recurrent event outcomes require careful thought for analysis, particularly when censoring of recurrent events is present. A method that can utilize many potentially highly correlated risk factors of recurrent events for dynamic prediction would be of interest to modelers and practitioners.
For dynamic prediction of an upcoming recurrent event, patient history of such events is often one of the most useful predictors (typically recurrent event-times within a patient are positively correlated). A feature of recurrent event data is that each patient’s event history at study time, $t$, becomes richer for making predictions as $t$ increases. For instance in the Azithromycin for the Prevention of Exacerbation of COPD study (Albert and others, 2011), recurrent exacerbations were monitored over a year of follow-up; a cohort that we would like to use for making dynamic predictions of exacerbations. Unfortunately, patient history data for the most recent exacerbation time prior to study enrollment was not collected in this study. As a result, this patient history data would not be available for dynamically predicting future events until later in study. How to include patient history data in dynamic predictions, when this data is only partially available at early follow-up times, has not been addressed in the recurrent-event setting.
@@
The majority of literature in recurrent event analysis has been parametric or semi-parametric in nature, with specific modeled relationships between risk factors and recurrent event outcomes. Models have been designed for the analysis of recurrent events, gap times, and times-to-first-event across different follow-up windows, each with different assumptions imposed by the modeling paradigms (e.g., Andersen and Gill, 1982; Miloslavsky and others, 2004; Ang and others, 2009; Mao and Lin, 2016; Hengelbrock and others, 2016; Kong and others, 2016; Chan and Wang, 2017a; Liang and others, 2017; Xu and others, 2017; Choi and others, 2017; Chan and Wang, 2017b; Xia, Murray and Tayob, 2020; Safari and others, 2023). Such models have interpretable parameter estimates that can inform medical literature, generate hypotheses, and to some degree, predict outcomes. Disadvantages include limits to estimating these same parameters when there are many predictors or collinearity between important predictors that make unique solutions for corresponding parameter estimates difficult to obtain (often due to difficulty with inversion of nearly singular matrices; see, for example, Hoerl and Kennard (1970); Gunst and Webster (1975)). In addition, methods that incorrectly assume independence of gap times between recurrent events measured on an individual give biased results due to induced dependent censoring.
A few authors propose pseudo-observation approaches to parametrically or semi-parametrically model censored recurrent event data (Yokota and Matsuyama, 2019; Xia, Murray and Tayob, 2020), replacing censored outcomes with tailor-made pseudo-observations suitable for obtaining desired parameter estimates from models available for uncensored data. Yokota and Matsuyama (2019) developed pseudo-observations for modeling probabilities that the $k$-th recurrent event occurred in a particular follow-up window. Xia, Murray and Tayob (2020) developed pseudo-observations for modeling the restricted mean event-free time in a particular follow-up window. Both of these approaches are versions of landmark analysis, where outcomes for analysis are measured relative to landmark times. Xia, Murray and Tayob (2020)’s landmarking approach transforms censored recurrent event data into a censored longitudinal data structure with restricted mean pseudo-observations defined over a series of follow-up windows that can be analyzed using standard longitudinal data analysis methods.
@@
As to the goal of dynamically predicting recurrent event risk based on traditional censored recurrent event data with many potentially colinear predictors, perhaps with non-linear associations with recurrent event outcomes, we are drawn to the nonparametric nature of random forest algorithms. Machine learning methods, and in particular random forest algorithms, have gained widespread acceptance as a method of building predictions based on complex relationships of biomarkers (first proposed by Breiman and others (1984), applied more recently in statistical literature such as Zhao and others (2020), Gerber and others (2021), and reviewed in Hu and Szymczak (2023)). These algorithms naturally accommodate non-canonical relationships between covariates and outcomes, an attractive feature in an age of increasingly high-dimensional data where a traditional modeler cannot realistically find all the interactions, higher-order covariates and non-linear patterns that may be predictive of an outcome. By using information across different sets of covariates within each of the generated trees of a random forest, a modeler can side-step several traditional pitfalls of parametric and semi-parametric models: (1) multicolinearity of predictors and (2) limited ability to include a large number of covariates. In particular, Zhao and others (2020) showed that for a single event time, imposing a censored longitudinal data structure introduced by Tayob and Murray (2015) and incorporating ideas from pseudo-observation literature, random forest regression can address nonlinear relationships between covariates and outcomes for dynamic prediction of event-free periods. Machine learning methods for censored recurrent event data are still in their infancy; we are not aware of any machine learning methods available for the analysis of recurrent event data.
In our manuscript we incorporate ideas from Xia, Murray and Tayob (2020) on building a censored longitudinal structure for censored recurrent event data using pseudo-observations. Our methods make several important advances from this point: (1) To better reflect the goals of dynamic prediction, we develop pseudo-observations for analysis of event-free probabilities over $\tau$-duration follow-up windows measured from regularly spaced landmark times over the duration of the entire follow-up period. (2) We develop a modified version of the Xia, Murray and Tayob (XMT) model where the estimand of interest is the probability of being event-free over the subsequent $\tau$-duration window, rather than a $\tau$-restricted mean. (3) We propose and evaluate different methods for including historical event rate information in these analyses with a multiple imputation approach for addressing missing event history data prior to the initiation of follow-up, and finally, (4), we describe a novel random forest algorithm with increased prediction accuracy to perform dynamic prediction of event-free probabilities over $\tau$-duration follow-up windows applicable to censored recurrent event data using elements from advances (1) and (3). This final random forest tool addresses many difficult issues facing analysts of recurrent event data mentioned above and we compare the performance of dynamic prediction using our modified-XMT model versus our random forest algorithm.
@@
The rest of this paper is organized as follows. In Section 2, we describe how to transform censored recurrent event data to a censored longitudinal data framework for analysis and offer examples of useful patient history covariates. In Section 3, we describe how to construct pseudo-observations based on the censored longitudinal data structure given in Section 2. Using these pseudo-observations, we describe dynamic prediction based on our modified version of the Xia, Murray and Tayob (XMT) model in Section 4. In Section 5, again using pseudo-observations defined in Section 3, we describe how to perform dynamic prediction via a random forest algorithm applicable to longitudinal data (i.e., the constructed pseudo-observations). Model evaluation metrics are summarized in Section 6. In Section 7, we use simulations to compare our modified XMT regression approach and our proposed random forest approach for the analysis of censored recurrent event data, and evaluate different methods for including historical event rate information in these analyses. We apply both the modified XMT and our recurrent event random forest algorithm to data from the study of Azithromycin for the Prevention of Exacerbation of COPD study in Section 8. We conclude with a discussion in Section 9.
@@ 2 Notation and Censored Longitudinal Data Structure
For individual $i$, denote recurrent event times $T_{i,1}<T_{i,2}<\ldots$, that are potentially censored at time $C_{i}$, where $C_{i}$ is independent of $T_{i,j}$, $i=1,...,n$, and $j\geq 1$. Independent censoring mechanisms are common, for instance, in settings where participants are identified during a finite accrual period and followed for a predetermined period (often based on research funding). Observed data for each individual is $X^{*}_{i,j}=\min(T_{i,j},C_{i})$, $i=1,\ldots,n$, and $j\geq 1$. These event times are sometimes converted to gap times between recurrent events $G_{i,j}=T_{i,j+1}-T_{i,j}$, $i=1,\ldots,n$, $j\geq 1$ for analyses. However, a well-known challenge is the dependence between gap time random variables, $G_{i,j}$, and corresponding censoring times, $C_{i}-T_{i,j-1}$ (Zhu, 2014).
One approach to avoiding dependent censoring in analyses of recurrent events was proposed by Tayob and Murray (2015), who transformed correlated recurrent event times into a censored longitudinal data structure. First, a series of follow-up window start times are pre-specified to $\mathcal{T}=\{t_{0},t_{1},t_{2},\ldots\}$, with $t_{j}-t_{j-1}=a$, $j\geq 1$; the start times can be as close as $a=1$ day apart, if desired. Computational speed of the analysis is usually the limiting factor in choosing $a$. Xia and Murray (2019) found that $a$ equal to one-third of the mean gap time in the control group resulted in capturing 90% of the recurrent events from the original data set in at least one follow-up window. In practice, more frequent follow-up window start times have the potential for more precise estimation, but with diminishing returns once all recurrent events have been captured in at least one follow up window.
In this paper, at each of the pre-specified “check-in time points” in $\mathcal{T}$, we will focus on estimating the probability of remaining recurrent-event-free over a subsequent follow-up window of length $\tau$, where $\tau$ is user-specified and ideally informed by clinical expertise. Although the check-in time points are shared across subjects by design, potential censoring may result in different numbers of effective check-ins; we therefore denote the subject-specific sets of check-in time points by $\{\mathcal{T}_{i}\}_{i=1}^{n}$, where $\mathcal{T}_{i}=\{t_{0},t_{1},...t_{b_{i}}\}$. In the case of no censoring, we chose $t_{b_{i}}$ so that the final follow-up interval from $t_{b_{i}}$ to $(t_{b_{i}}+\tau)$ does not exceed the study period; in the case with censoring, we choose $t_{b_{i}}<C_{i}$. We follow Tayob and Murray (2015) and construct the censored longitudinal data that is essential for creating pseudo-observations in Section 3. We repeat the following two steps at all the check-in time points $t\in\mathcal{T}_{i}$, $i=1,\ldots,n$:
@@
    Identify the first recurrent event occurred after $t$; let $\eta_{i}(t)=\min\{j:X^{*}_{i,j}\geq t,j=1,\ldots,n_{i}\}$ be its index among all the observed recurrent events for subject $i$. Define the residual recurrent-event-free time by $T_{i}(t)=T_{i,\eta_{i}(t)}-t$ with the corresponding censoring random variable $C_{i}(t)=C_{i}-t$, which importantly remains independent of $T_{i}(t)$;
@@
    Construct the observed censored longitudinal data by $\{X_{i}(t)=X^{*}_{i,\eta_{i}(t)}-t,\delta_{i}(t)=I\{X_{i}(t)\leq C_{i}(t)\}:i=1,2,\ldots,n\}$.